find set of a which equation x(x+1)(x+a(x+a+1)=a² has four real roots

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear piyush

x(x+1)(x+a)(x+a+1)=a²

x(x+1)(x+a){(x+1)+a}=a²

x(x+1){x(x+1)+2(xa)+a +a2}=a²

{x(x+1)}2 + 2(xa){x(x+1)} + (a+a2) x(x+1) =a2

{x(x+1)}2 + 2(xa){x(x+1)} + (ax)2  +ax(x+1+a)  =a2

 {x(x+1) +ax}2 + a{x(x+1) +ax} =a2

 {x(x+1) +ax +a/2}2   =(√5a/2)2

{x(x+1) +ax +a/2 -√5a/2}{x(x+1) +ax +a/2 +√5a/2} =0

{x2 +x(a+1) +a/2 -√5a/2}{x2 +x(a+1) +a/2 +√5a/2} =0

x2 +x(a+1) +a/2 +√5a/2 =0   and  x2 +x(a+1) +a/2 -√5a/2 =0

for 4 real value discrimnent of both the equation must be grater than or equal to zero

x2 +x(a+1) +a/2 +√5a/2 =0            and       x2 +x(a+1) +a/2 -√5a/2 =0

D≥0                                                                      D≥0

(a+1)2 -4(a/2 +√5a/2)≥0                      (a+1)2 -4(a/2 -√5a/2)≥0                 

  (a-(√5+2))(a-(√5-2))≥0                      (a-(-√5+2))(a-(-√5-2))≥0

 so solution is

(-∞,-2-√5]U[2-√5,-2+√5]U[2+√5,∞)

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin